# Geometric control of boundary-catalytic branching processes

### v1

@@ Abstract
and the principal eigenvalue $\mu_{0}$ gives the critical catalytic rate $q_{\mathrm{c}}^{\rm crit}$ (here the eigenpairs are denoted as $\{\mu,v\}$ to distinguish them from the previous problem). Moreover, we show (see SM.IV) that the associated eigenfunction $v_{0}$ determines the asymptotic behavior of $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$ as $q_{\mathrm{c}}$ approaches $q_{\mathrm{c}}^{\rm crit}$:
@@
which agrees with Eq. (7) for a spherical shell. The identification of this general mathematical framework for relating the parameters $q_{\mathrm{a}}$, $q_{\mathrm{c}}$ and $\lambda_{0}$ and for determining $q_{\mathrm{c}}^{\rm crit}$ opens promising perspectives for geometric control using spectral geometry and related asymptotic and numerical tools. For instance, Fig. 2(c) shows the optimal absorption rate $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$ for two configurations (illustrated on panels 2(a,b)) with one catalytic region and nine randomly located absorbing regions of circular shape. In the low-rate regime, the approximate relation (5) is valid and provides the lower bound on $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$, as expected. In turn, as $q_{\mathrm{c}}$ approaches the critical value $q_{\mathrm{c}}^{\rm crit}$ (shown by vertical dashed lines), $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$ diverges according to Eq. (10). Figure 2(c) can also be interpreted as the phase diagram in the space of $q_{\mathrm{c}}$ and $q_{\mathrm{a}}$ rates, in which the optimal absorption rate $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$ is the critical line that separates an exponential growth of population from its extinction. Moreover, when the catalytic and absorbing regions are small, matched asymptotic expansions can be applied to estimate the critical value $q_{\mathrm{c}}^{\rm crit}$ without solving the spectral problem (9) (see SM.V for details).
@@
Figure 2: (a,b) Two configurations composed of one catalytic circle of radius $R=0.1$ at the center, and 9 randomly located absorbing circles of radius $R_{\mathrm{a}}$ (with $R_{\mathrm{a}}=0.05$ (a) and $R_{\mathrm{a}}=0.1$ (b)), enclosed by a reflecting circle of radius $L=1$. Colors ranging from dark blue to dark red represent variations of the limiting population size $N(\infty|\bm{x})$ for the special case $q_{\mathrm{a}}=\infty$ (perfect sinks) and $q_{\mathrm{c}}=q_{\mathrm{c}}^{\rm crit}$. (c) The optimal absorption rate $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$ as a function of the catalytic rate $q_{\mathrm{c}}$ for two considered configurations (shown by solid line for $R_{\mathrm{a}}=0.05$ and by dashed line for $R_{\mathrm{a}}=0.1$). Dotted lines show the low-rate approximation (5), whereas dash-dotted lines present the asymptotic relation (10), with $Q\approx 15.78$ for $R_{a}=0.05$ and $Q\approx 19.21$ for $R_{a}=0.1$. Vertical lines indicate the critical values $q_{\mathrm{c}}^{\rm crit}\approx 6.61$ and $q_{\mathrm{c}}^{\rm crit}\approx 7.62$ for two configurations. These numerical results were obtained by a finite-element method [Chaigneau24], with the maximal meshsize $h_{\rm max}=0.01$ (see SM.VI for details).
@@
Up to this point, we focused on bounded domains. The explicit description (7, 8) of the steady-state solution in a spherical shell helps to clarify what happens for unbounded domains when the absorbing region $\Gamma_{\mathrm{a}}$ is (re)moved to infinity. Looking at the limit of Eq. (8) as $L\to\infty$, one can appreciate the crucial difference between two-dimensional and higher-dimensional settings. In two dimensions, the critical value $q_{\mathrm{c}}^{\rm crit}$ logarithmically vanishes as $L\to\infty$. In other words, if the absorbing boundary $\Gamma_{\mathrm{a}}$ is pushed to infinity, one gets $q_{\mathrm{c}}^{\rm crit}=0$ so that, for any $q_{\mathrm{c}}>0$, the population grows exponentially. This is a consequence of the recurrent nature of Brownian motion in the plane: each particle never stops returning to the catalytic surface that surely triggers its branching. The situation is different in higher dimensions, where $q_{\mathrm{c}}^{\rm crit}\to(d-2)/R$ as $L\to\infty$. Even though the absorbing region is removed, eventual escapes of particles to infinity due to the transient character of Brownian motion substitute the absorption events. According to the explicit computation of $N(t|\bm{x})$ for the exterior of a three-dimensional ball of radius $R$ (see SM.VII for details), one can distinguish three scenarios: (i) if $q_{\mathrm{c}}>1/R$, the branching events are more efficient than the escape events, so that the population size $N(t|\bm{x})$ grows exponentially, despite the unboundedness of the environment; (ii) if $q_{\mathrm{c}}<1/R$, the population size reaches a finite steady-state value $N(\infty|\bm{x})=1+(R/|\bm{x}|)/(1/(q_{\mathrm{c}}R)-1)$; (iii) in the critical setting $q_{\mathrm{c}}=1/R$, the population size exhibits an unlimited power-law growth: $N(t|\bm{x})\propto\sqrt{Dt}/|\bm{x}|$, as $t\to\infty$. Moreover, the critical value $1/R$ can be identified again with the principal eigenvalue $\mu_{0}$ of the Steklov problem (9) in the exterior of a ball, for which $\Gamma_{\mathrm{a}}=\Gamma_{\mathrm{r}}=\emptyset$.
@@
While the above picture was based on the explicit solution for the exterior of a ball, it remains valid for general exterior domains. In fact, let $\mu_{0}$ be the principal Steklov eigenvalue in the exterior $\Omega={\mathbb{R}}^{d}\backslash\overline{\Omega_{0}}$ of a bounded domain $\Omega_{0}$ with boundary $\partial\Omega=\Gamma_{\mathrm{c}}$ (see [Auchmuty14, Arendt15] for its rigorous formulation). The Laplace operator in $\Omega$ with Robin boundary condition on $\Gamma_{\mathrm{c}}$ has (at least) one strictly negative eigenvalue, if and only if $q_{\mathrm{c}}>q_{\mathrm{c}}^{\rm crit}=\mu_{0}$ [Bundrock25]. As a consequence, even though the spectral expansion (3) is not applicable because the Laplace operator has the essential spectrum $0,+\infty)$, the presence of the isolated eigenvalue $\lambda_{0}<0$ ensures an exponential growth of the population size with the rate $1/(D|\lambda_{0}|)$ (see [Freitas15, Krejcirik18, Krejcirik20] for further discussions of the related spectral properties). In contrast, the analysis of the long-time asymptotic behavior of the population size in the case $q_{c}\leq\mu_{0}$ remains an open problem.
@@
Throughout this Letter, we assumed that the particles could be killed exclusively on the absorbing region $\Gamma_{\mathrm{a}}$. In many applications, the bulk medium itself can be aggressive and lead to eventual destructions of the diffusing particles with some rate $\nu>0$ [Yuste13, Meerson15, Grebenkov17]. This mechanism, which can also account for a finite lifetime of particles and their spontaneous death in the bulk, can be easily incorporated by adding the reactive term $\nu N(t|\bm{x})$ to the right-hand side of the diffusion equation (1). In this case, the steady-state regime would correspond to $\lambda_{0}=-\nu/D$ instead of $\lambda_{0}=0$, and the above analysis can be adapted to this more general situation.

### v2

@@ III.4 Phase diagram for bounded domains
and the principal eigenvalue $\mu_{0}$ gives the critical catalytic rate $q_{\mathrm{c}}^{\rm crit}$ (here the eigenpairs are denoted as $\{\mu,v\}$ to distinguish them from the previous problem). Moreover, we show (see Appendix E) that the associated eigenfunction $v_{0}$ determines the asymptotic behavior of $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$ as $q_{\mathrm{c}}$ approaches $q_{\mathrm{c}}^{\rm crit}$:
@@
which agrees with Eq. (7) for a spherical shell. The identification of this general mathematical framework for relating the parameters $q_{\mathrm{a}}$, $q_{\mathrm{c}}$ and $\lambda_{0}$ and for determining $q_{\mathrm{c}}^{\rm crit}$ opens promising perspectives for the geometric control using spectral geometry and related asymptotic and numerical tools.
@@
Figure 2: (a,b) Two configurations composed of one catalytic circle of radius $R=0.1$ at the center, and 9 randomly located absorbing circles of radius $R_{\mathrm{a}}$ (with $R_{\mathrm{a}}=0.05$ (a) and $R_{\mathrm{a}}=0.1$ (b)), enclosed by a reflecting circle of radius $L=1$. Colors ranging from dark blue to dark red represent variations of the limiting mean population size $N(\infty|\bm{x})$ for the special case $q_{\mathrm{a}}=\infty$ (perfect sinks) and $q_{\mathrm{c}}=q_{\mathrm{c}}^{\rm crit}$. (c) The optimal absorption rate $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$ as a function of the catalytic rate $q_{\mathrm{c}}$ for two considered configurations (shown by solid line for $R_{\mathrm{a}}=0.05$ and by dashed line for $R_{\mathrm{a}}=0.1$). Dotted lines show the low-rate approximation (5), whereas dash-dotted lines present the asymptotic relation (10), with $Q\approx 15.78$ for $R_{a}=0.05$ and $Q\approx 19.21$ for $R_{a}=0.1$. Vertical lines indicate the critical values $q_{\mathrm{c}}^{\rm crit}\approx 6.61$ and $q_{\mathrm{c}}^{\rm crit}\approx 7.62$ for two configurations. These numerical results were obtained by a finite-element method [43], with the maximal meshsize $h_{\rm max}=0.01$ (see Appendix F for details).
Figure 2(c) shows the optimal absorption rate $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$ for two configurations (illustrated on panels 2(a,b)) with one catalytic region and nine randomly located absorbing regions of circular shape. This rate was computed numerically by solving the Steklov spectral problem via a finite-element method [43] (see Appendix F). In the low-rate regime, the approximate relation (5) is valid and provides the lower bound on $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$, as expected. In turn, as $q_{\mathrm{c}}$ approaches the critical value $q_{\mathrm{c}}^{\rm crit}$ (shown by vertical dashed lines), $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$ diverges according to Eq. (10). As said earlier, Fig. 2(c) can be interpreted as the phase diagram in the space of $q_{\mathrm{c}}$ and $q_{\mathrm{a}}$ rates, in which the optimal absorption rate $\hat{q}_{\mathrm{a}}(q_{\mathrm{c}},0)$ is the critical line that separates an exponential growth of population from its extinction.
III.5 Asymptotic results for small regions
When the catalytic and absorbing regions are small (Fig. 2(a,b)), matched asymptotic expansions can be applied to estimate the principal Steklov eigenvalue $\mu_{0}$ and thus the critical value $q_{\mathrm{c}}^{\rm crit}$ without solving the spectral problem (9). In two dimensions, the asymptotic analysis of the Steklov problem (9) was undertaken in [44] for the case when the reflecting boundary $\partial\Omega$ is smooth, connected and covered by small catalytic and absorbing regions. If the catalytic region is a single arc-like set, whereas the absorbing region is a finite union of well-separated arc-like sets, the critical catalytic rate is
$q_{\mathrm{c}}^{\rm crit}=\mu_{0}\simeq\frac{1}{\epsilon_{\rm c}\bigl[C\ln(L/\epsilon_{\rm c})+C_{0}\bigr]}\,$, (11)
where $2L$ is the diameter of the domain $\Omega$, $2\epsilon_{\rm c}$ is the perimeter of the catalytic region $\Gamma_{\rm c}$, $C_{0}=\frac{2}{\pi}\bigl(\frac{3}{2}-\ln 2\bigr)$ is a numerical constant, whereas $C$ depends on the lengths of all patches, their locations on the boundary $\partial\Omega$, and the shape of the confining domain (see Eq. (35) and Appendix G for details). For instance, if the domain $\Omega$ is the disk and the absorbing region is an arc of length $2\epsilon_{\rm a}$, the factor $C$ can be found explicitly, yielding
$q_{\mathrm{c}}^{\rm crit}\approx\frac{\pi}{2\epsilon_{\rm c}}\biggl(-\ln(\epsilon_{\rm c}\epsilon_{\rm a})+\frac{3}{2}+2\ln|\bm{x}_{\rm c}-\bm{x}_{\rm a}|\biggr)^{-1},$ (12)
where $\bm{x}_{\rm c}$ and $\bm{x}_{\rm a}$ are the positions of the catalytic and absorbing regions on the boundary. Expectedly, when the catalytic region $\Gamma_{\mathrm{c}}$ is reduced (i.e., $\epsilon_{\rm c}\to 0$), the critical catalytic rate $q_{\mathrm{c}}^{\rm crit}$ increases, so that it is easier to compensate the effect of branching events. In contrast, if the absorbing region $\Gamma_{\mathrm{a}}$ is reduced (i.e., $\epsilon_{\rm a}\to 0$), less particles can be trapped, and $q_{\mathrm{c}}^{\rm crit}$ decreases. Similarly, the critical catalytic rate also decreases when the distance $|\bm{x}_{\mathrm{c}}-\bm{x}_{\mathrm{a}}|$ between two regions increases. However, both effects are weak due to the logarithmic nature of the terms $\ln(1/\epsilon_{\rm a})$ and $\ln|\bm{x}_{\mathrm{c}}-\bm{x}_{\mathrm{a}}|$. In Appendix G, we provide an extension of this asymptotic analysis to the case when the catalytic region is a small disk inside $\Omega$, whereas the absorbing region is a finite union of small well-separated disks, like the configurations shown in Fig. 2(a,b). A similar asymptotic analysis of the Steklov problem (9) in three dimensions is provided [45].
III.6 Extension to unbounded domains
@@
Up to this point, we focused on bounded domains. The explicit description (7, 8) of the steady-state solution in a spherical shell helps to clarify what happens for unbounded domains when the absorbing region $\Gamma_{\mathrm{a}}$ is (re)moved to infinity. Looking at the limit of Eq. (8) as $L\to\infty$, one can appreciate the crucial difference between two-dimensional and higher-dimensional settings. In two dimensions, the critical value $q_{\mathrm{c}}^{\rm crit}$ logarithmically vanishes as $L\to\infty$. In other words, if the absorbing boundary $\Gamma_{\mathrm{a}}$ is pushed to infinity, one gets $q_{\mathrm{c}}^{\rm crit}=0$ so that, for any $q_{\mathrm{c}}>0$, the population grows exponentially. This is a consequence of the recurrent nature of Brownian motion in the plane: each particle never stops returning to the catalytic surface that surely triggers its branching. The situation is different in higher dimensions, where $q_{\mathrm{c}}^{\rm crit}\to(d-2)/R$ as $L\to\infty$. Even though the absorbing region is removed, eventual escapes of particles to infinity due to the transient character of Brownian motion substitute the absorption events. According to the explicit computation of $N(t|\bm{x})$ for the exterior of a three-dimensional ball of radius $R$ (see Appendix H for details), one can distinguish three scenarios: (i) if $q_{\mathrm{c}}>1/R$, the branching events are more efficient than the escape events, so that the mean population size $N(t|\bm{x})$ grows exponentially, despite the unboundedness of the environment; (ii) if $q_{\mathrm{c}}<1/R$, the mean population size reaches a finite steady-state value $N(\infty|\bm{x})=1+(R/|\bm{x}|)/(1/(q_{\mathrm{c}}R)-1)$; (iii) in the critical setting $q_{\mathrm{c}}=1/R$, the mean population size exhibits an unlimited power-law growth: $N(t|\bm{x})\propto\sqrt{Dt}/|\bm{x}|$, as $t\to\infty$. Moreover, the critical value $1/R$ can be identified again with the principal eigenvalue $\mu_{0}$ of the Steklov problem (9) in the exterior of a ball, for which $\Gamma_{\mathrm{a}}=\Gamma_{\mathrm{r}}=\emptyset$.
@@
While the above picture was based on the explicit solution for the exterior of a ball, it remains valid for general exterior domains. In fact, let $\mu_{0}$ be the principal Steklov eigenvalue in the exterior $\Omega={\mathbb{R}}^{d}\backslash\overline{\Omega_{0}}$ of a bounded domain $\Omega_{0}$ with boundary $\partial\Omega=\Gamma_{\mathrm{c}}$ (see [46, 47] for its rigorous formulation). The Laplace operator in $\Omega$ with Robin boundary condition on $\Gamma_{\mathrm{c}}$ has (at least) one strictly negative eigenvalue, if and only if $q_{\mathrm{c}}>q_{\mathrm{c}}^{\rm crit}=\mu_{0}$ [48]. As a consequence, even though the spectral expansion (3) is not applicable because the Laplace operator has the essential spectrum $0,+\infty)$, the presence of the isolated eigenvalue $\lambda_{0}<0$ ensures an exponential growth of the mean population size with the rate $1/(D|\lambda_{0}|)$ (see [49, 50, 51] for further discussions of the related spectral properties). In contrast, the analysis of the long-time asymptotic behavior of the mean population size in the case $q_{c}\leq\mu_{0}$ remains an open problem.
@@
Throughout this paper, we assumed that the particles could be killed exclusively on the absorbing region $\Gamma_{\mathrm{a}}$. In many applications, the bulk medium itself can be aggressive and lead to eventual destructions of the diffusing particles with some rate $\nu>0$ [52, 53, 54]. This mechanism, which can also account for a finite random lifetime of particles and their spontaneous death in the bulk, can be easily incorporated by adding the reactive term $\nu N(t|\bm{x})$ to the right-hand side of the diffusion equation (1). In this case, the steady-state regime would correspond to $\lambda_{0}=-\nu/D$ instead of $\lambda_{0}=0$, and the above analysis can be adapted to this more general situation.
IV Conclusions